# Review of Demographic Bias in Face Recognition

### v1

@@ Abstract
Demographic bias in face recognition (FR) has emerged as a critical area of research, given its impact on fairness, equity, and reliability across diverse applications. As FR technologies are increasingly deployed globally, disparities in performance across demographic groups– such as race, ethnicity, and gender– have garnered significant attention. These biases not only compromise the credibility of FR systems but also raise ethical concerns, especially when these technologies are employed in sensitive domains. This review consolidates extensive research efforts providing a comprehensive overview of the multifaceted aspects of demographic bias in FR.
@@ I Introduction
Demographic bias in face recognition (FR) systems has emerged as a critical challenge in the deployment of biometric technologies for real-world applications [1, 2, 3, 4]. Bias in these systems often leads to disparities in performance across demographic groups- such as variations in recognition accuracy- based on race, gender, and age [4, 5, 6]. Such biases can have far-reaching consequences, especially in critical applications like border crossing, law enforcement [7, 8], security [9], and hiring processes [10, 11, 12], where fairness and accuracy are paramount. Fig. 1 illustrates the issue of demographic bias based on race, gender, and age. Ideally, a fair model should exhibit equitable performance across all demographic groups (i.e. similar error rates or comparable score distributions). The disparity between groups contributes to the persistence of demographic bias. This issue is further compounded by the growing reliance on FR technologies, making it imperative to identify, evaluate, and mitigate sources of bias effectively.
Figure 1: Schematic illustration of issue of demographic bias in FR with reference to different demographic factors: (a) race or ethnicity, (b) gender, and (c) age.
@@
Due to its severity and widespread range of applications, demographic bias has emerged as a crucial area of research, drawing significant attention from both the biometrics and computer vision communities [3, 4, 13, 14]. This issue has been formally incorporated into the evaluation frameworks of prominent initiatives, such as the National Institute of Standards and Technology’s (NIST) Face Recognition Vendor Tests (FRVT), which have included demographic effects in their reports since 2019 [15, 16]. Similarly, the Maryland Test Facility (MdTF), supported by the United States Department of Homeland Security (DHS), has conducted biometric technology rallies to evaluate demographic disparities in FR systems [^1] MdTF. In Europe, organizations like the European Association for Biometrics (EAB) have hosted dedicated events on demographic fairness in biometric systems, underlining the global importance of this topic [^2] EAB event on Demographic Fairness in Biometric Systems. This research area is often positioned within the broader context of fairness and trustworthy biometrics, and has been receiving substantial attention– in the form of papers, workshops, or special sessions– from leading conferences such as IEEE/CVF CVPR, WACV, IEEE FG, and ICPR, and from reputable journals including IEEE Transactions on information forensics and security (TIFS), IEEE Transactions on Biometrics, Behavior, and Identity Science (TBIOM) [17] and IEEE Signal Processing Magazine [18]. Additionally, standards organizations have recognized the need for systematic approaches, with ISO/IEC 19795-10 recently publishing guidelines for quantifying demographic differentials in biometric systems emphasizing the need for addressing bias in FR technologies [19].
@@
Scope: Fairness and bias in machine learning are expansive topics, and their application in biometrics has drawn significant attention in recent years. Several comprehensive reviews have addressed fairness and bias in machine learning broadly [5, 20], while others focus specifically on biometrics, offering insights into various modalities such as face, fingerprint, and vein, alongside applications beyond recognition, including region of interest (ROI) detection, quality assessment, and presentation attack detection[3, 4, 13, 21]. However, as face remains the most commonly used biometric trait, a substantial portion of research on demographic bias has concentrated on this modality.
@@
While previous reviews offer broad perspectives, the extensive literature and emerging challenges specific to demographic bias in FR necessitate a dedicated review. In this article, we provide a consolidated discussion of recent advancements in the field, addressing the causes of demographic bias, available datasets for research, evaluation metrics for assessment of bias, and recent mitigation techniques. Furthermore, we explore ongoing challenges that persist in addressing demographic disparities, particularly in the light of novel use cases and emerging FR applications. Although our primary focus is on race and ethnicity, as these are dominant areas of research, we also include gender-related studies within the broader context. Age-related studies, given their distinct nature and established body of research, are referenced only where directly relevant.
@@
Naming Conventions: In this review, we adopt the terminologies used in the referenced works while acknowledging the minor differences in naming conventions. Terms such as race and ethnicity are often used interchangeably, although they represent distinct concepts. For clarity, we retain the terms employed by the original studies. Similarly, the names of ethnic groups vary across the literature: for instance, some works use terms like Black and White, while others prefer African and Caucasian. Additionally, the South Asian group is sometimes referred to as Indian, whereas Asians often refer to East Asians. To maintain consistency and respect the source material, we adhere to the original terminologies in this review. Readers are encouraged to refer to the cited works for precise definitions and context. While fairness in related topics such as face detection, image quality, expression recognition, and attribute estimation is of interest to the research community, this review exclusively focuses on demographic bias in FR.
@@
Contributions: This review constitutes the first comprehensive work dedicated to exploring demographic bias in FR, offering a unified and holistic perspective to researchers in the field. We systematically analyze and organize key aspects, including the causes of demographic bias, available datasets, assessment metrics, and mitigation techniques, providing a structured framework for understanding these areas. Finally, we identify emerging challenges and unresolved questions, inviting further research and innovation to advance fairness in FR systems.
@@
The structure of this paper is as follows: Section II explores the causes of bias, analyzing factors such as distribution of demographic groups in dataset, skin-tone, image quality, and algorithmic sensitivities. We provide an overview of datasets commonly used for bias-related research in Section III, highlighting their demographic attributes and suitability for specific tasks. Section IV reviews existing metrics for bias evaluation, discussing their strengths and limitations. Section V outlines recent bias mitigation strategies across different stages of the FR pipeline. We discuss open challenges and future research directions in Section VI, and conclude the review in Section VII.
II Causes of Demographic Bias
@@
In this section, we consolidate findings from existing works related to causes of demographic bias in FR. Considering the wide range of research in this area, we have grouped the causes into categories for clarity and systematic review. These categories encompass factors such as imbalances in training datasets, variability in skin tones, algorithmic sensitivity, image quality and related covariates, as well as combined or intersectional demographic factors. While this categorization simplifies the organization, it is important to note that many studies attribute demographic bias to multiple, overlapping factors, making strict classification difficult. We have categorized works based on their primary focus or findings, but some studies have been referenced across multiple categories to reflect their broader relevance. This approach also highlights that causes of demographic bias are inherently multifaceted and interconnected, requiring a thorough understanding to analyze and mitigate these disparities. A quick summary of various works discussed in this section can be found in Table I.
@@ II-A Training Datasets
The issue of demographic bias in FR systems often stems from imbalanced or unrepresentative datasets, significantly influencing both training and evaluation outcomes. Research by Krishnapriya et al. [22] demonstrated how demographic groups, such as African-American cohorts, exhibit higher false match rates (FMR), while Caucasian cohorts face higher false non-match rates (FNMR), highlighting the interplay between ethnicity and matching thresholds. The Face Recognition Vendor Test (FRVT) conducted by NIST [15] substantiated these findings, reporting increased false positives in women, children, and the elderly, alongside higher false negatives in under-represented racial groups, emphasizing the intricate interactions between dataset characteristics and demographic attributes.
@@
Early studies like Klare et al.  [23] advocated for balanced datasets and the use of exclusive cohorts to enhance FR performance. Cavazos et al. [24] identified how dataset complexity and identification thresholds contribute to racial bias, such as the need for higher thresholds for East-Asian faces to achieve comparable false acceptance rates (FARs). Gwilliam et al. [25] challenged the prevailing assumptions about the necessity of balanced datasets by demonstrating that skewed distributions favoring African faces reduced racial bias more effectively than balanced datasets. Wu and Bowyer [26] expanded this discussion, emphasizing that mere balance in identities or number of images is insufficient to address bias, highlighting additional factors like brightness and head pose during dataset assembly.
@@
Other works delved into specific aspects of demographic balance in datasets. Wang et al.  [27] observed that even race-balanced datasets failed to eliminate racial bias, hypothesizing that certain ethnicities are inherently more challenging to recognize. Kolla and Savadamuthu [28] highlighted the influence of facial quality and racial feature gradations on model fairness. Focusing on inter-sectional bias, Muthukumar et al.  [29] identified structural facial features as significant contributors, particularly for dark-skinned females, over attributes like skin tone or hair length. Cook et al. [30] further analyzed the role of image acquisition conditions, noting how factors such as skin reflectance and environmental conditions disproportionately affect darker-skinned individuals, thus advocating for standardized acquisition protocols to mitigate bias.
Although most studies acknowledge that demographically imbalanced training data contribute to biased FR models and training with balanced datasets enhances fairness, there is a consensus that these are neither the sole causes nor complete solutions to the broader issue of demographic bias.
@@ II-B Variability in Skin-tone
The influence of skin tone on the performance of FR systems has been extensively studied, revealing significant demographic disparities. In [29], Muthukumar et al. identified notable under-performance in recognizing dark-skinned females compared to other demographic group for commercial classifiers. Their analysis attributed these disparities to structural features such as lips, eyes, and cheeks, in addition to skin-tone itself. Similarly, Buolamwini and Gebru [31] employed the Fitzpatrick skin classification system to evaluate gender classifiers and reported the lowest accuracy for darker-skinned females. Their findings further indicated that lighter-skinned males achieved the highest performance, highlighting the intersection of skin-tone and gender as critical factors influencing recognition accuracy. In another study, Krishnapriya et al.  [22] examined FMR and FNMR across skin-tone groups, observing higher FMR for African-American cohorts and higher FNMR for Caucasians; however, they did not find a direct causation between darker skin tones and higher error rates.
@@
The Biometric Technology Rallies organized by MdTF have offered comprehensive insights into the role of skin-related factors in FR bias. Their 2019 report [30] emphasized skin reflectance as more significant predictor of performance disparities than race. Using systematic linear modeling, their study demonstrated that darker skin-tones were associated with longer transaction (processing overall pipeline) times and lower accuracy in biometric systems. This dependency was found to vary substantially across systems, highlighting important role of acquisition methods in determining the extent of bias. Lu et al.  [32] provided a quantitative assessment of performance variations across five skin-tone groups, identifying light-skinned individuals as the easiest to verify and darker-skinned individuals as the most challenging. However, ambiguities in defining skin tone categories complicate direct evaluations, highlighting the need for standardized classification metrics.
II-C Algorithmic Factors
In this section, we examine the sensitivity and limitations of algorithms in addressing demographic attributes. Phillips et al. [33] identified the “other-race effect,” where algorithms developed in Western and East Asian contexts demonstrated superior performance for their respective majority racial groups. This disparity persisted even when datasets were balanced, pointing to underlying biases in algorithmic design and training processes. Klare et al.  [23] observed recognition challenges for specific demographic groups, including females, Black individuals, and younger cohorts. They reported improved performance when models were trained exclusively on these groups, emphasizing the impact of training data composition. In [34], Nagpal et al. demonstrated that deep learning models encode in-group biases, mirroring human tendencies such as own-race and own-age effects. By analyzing activation maps, they showed that these biases were ingrained within the feature representations of the models. Wang et al. [27], using the Racial Faces in-the-Wild (RFW) dataset, validated racial bias, revealing that error rates for African faces were nearly double those for Caucasians, even with race-balanced training data. Serna et al. [35] similarly highlighted significant performance gaps across demographic groups, advocating for diverse training datasets and fairness-aware algorithmic designs.
Further investigations into gender-based biases by Albiero et al. [36] revealed skewed impostor and genuine score distributions as the primary reasons for lower accuracy in women. This bias persisted across datasets, regardless of balanced training and neutral facial expressions. Ricanek et al. [37] noted unique challenges in recognizing children’s faces due to structural changes with age, finding that algorithms effective on adult faces performed poorly for younger subjects.
@@ II-C Algorithmic Factors
Buolamwini and Gebru [31] 2018 IJB-A, Adience GN, ST Demonstrated lowest classifier performance for darker-skinned females using the Fitzpatrick system.
@@
Klare et al. [23] 2012 PCSO ET, GN, AG Highlighted recognition challenges for female, Black, and younger cohorts, improved with exclusive group training.
@@
Ricanek et al. [37] 2015 ITWCC AG Observed challenges in recognizing children’s faces due to structural changes with age, affecting algorithm accuracy.
@@
Sarridis et al. [44] 2023 RFW ET, GN, AG Reported high mistreatment rates for African females over 60 years, highlighting compounded biases.
@@
TABLE I: Summary of works delving into various causes of demographic bias in face recognition. As several works have identified multiple causes of bias, we have categorized the works based on their primary focus or inference. For details, readers are encouraged to refer to the source materials. The demographic factors of primary interest are denoted as ET: Ethnicity or race, GN: gender or sex, AG: age; whereas + indicates study of more attributes.

### v2

# Review of Demographic Fairness in Face Recognition

@@ Abstract
Demographic fairness in face recognition (FR) has emerged as a critical area of research, given its impact on fairness, equity, and reliability across diverse applications. As FR technologies are increasingly deployed globally, disparities in performance across demographic groups– such as race, ethnicity, and gender– have garnered significant attention. These biases not only compromise the credibility of FR systems but also raise ethical concerns, especially when these technologies are employed in sensitive domains. This review consolidates extensive research efforts providing a comprehensive overview of the multifaceted aspects of demographic fairness in FR.
@@ I Introduction
Demographic fairness in face recognition (FR) systems has emerged as a critical challenge in the deployment of biometric technologies for real-world applications [1, 2, 3, 4]. Disparities in these systems often leads to disparities in performance across demographic groups- such as variations in recognition accuracy- based on race, gender, and age [4, 5, 6]. Such disparities, also commonly referred to as biases, can have far-reaching consequences, especially in critical applications like border crossing, law enforcement [7, 8], security [9], and hiring processes [10, 11, 12], where fairness and accuracy are paramount.
A conventional FR system evaluates the similarity between two facial images by computing a matching score based on their extracted features to determine whether they correspond to the same individual or not. The performance of such systems is typically examined by analyzing the distributions of matching scores for image pairs of the same identity (referred to as mated pairs) and those of different identities (non-mated pairs). Scores are thresholded into binarized decisions: match or no-match. Fig. 1 illustrates the issue of demographic bias based on race, gender, and age. Ideally, a fair FR model should exhibit equitable performance across all demographic groups (i.e., similar error rates or comparable score distributions). The disparity between groups contributes to the unfair or biased FR. (The formal definitions for error rates and disparities provided in Sec. IV.) As depicted by various choices of scoring thresholds, accuracy differentials exist across demographic splits and can manifest at various thresholds. This emphasizes that demographic disparities are not tied solely to false match or false non-match errors but may be present throughout the score space. This issue is further compounded by the growing reliance on FR technologies, making it imperative to identify, evaluate, and mitigate sources of accuracy differentials effectively.
Figure 1: Illustration of differential performance across demographic groups in FR with reference to different demographic factors: (a) race or ethnicity, (b) gender, and (c) age. Score distributions for mated and non-mated pairs are shown for multiple demographic groups, along with decision thresholds (dotted lines) corresponding to different operating points. This figure is intended to demonstrate that accuracy differentials may persist across a range of thresholds and are not restricted to a particular type of error (e.g., false match or false non-match).
@@
Due to its severity and the widespread range of FR applications, demographic fairness has emerged as a crucial area of research, drawing significant attention from both the biometrics and computer vision communities [3, 4, 13, 14]. This issue has been formally incorporated into the evaluation frameworks of prominent initiatives, such as the National Institute of Standards and Technology’s (NIST) Face Recognition Vendor Tests (FRVT), which have included demographic effects in their reports since 2019 [15, 16]. Other initiatives, including the Maryland Test Facility (MdTF) and the European Association for Biometrics (EAB), have also contributed to this discourse, though at a smaller scale and scope compared to NIST. The MdTF, supported by the United States Department of Homeland Security (DHS), has conducted biometric technology rallies to evaluate demographic disparities in FR systems. In Europe, organizations such as the EAB have hosted dedicated events on demographic fairness in biometric systems, underlining the global importance of this topic. This research area is often positioned within the broader context of fair and trustworthy biometrics, and has been receiving substantial attention– in the form of papers, workshops, or special sessions– from leading conferences such as IEEE/CVF CVPR, WACV, IEEE FG, and ICPR, and from reputable journals including IEEE Transactions on Information Forensics and Security (TIFS), IEEE Transactions on Biometrics, Behavior, and Identity Science (TBIOM) [17] and IEEE Signal Processing Magazine [18]. Additionally, standards organizations have recognized the need for systematic approaches. For example, the recently published ISO/IEC 19795-10 guidelines quantify demographic differentials in biometric systems, and emphasize the need for addressing fairness in FR technologies [19].
@@
Scope: Fairness and bias in machine learning are expansive topics, and their application in biometrics has drawn significant attention in recent years. Several comprehensive reviews have addressed fairness and bias in machine learning broadly [5, 20], while others focus specifically on biometrics, offering insights into various modalities such as face, fingerprint, and vein, alongside applications beyond recognition, including region of interest (ROI) detection, quality assessment, and presentation attack detection[3, 4, 13, 21]. However, as face remains the most commonly used biometric trait in operational scenarios, a substantial portion of research on demographic bias has concentrated on this modality.
@@
While previous reviews offer broad perspectives, the extensive literature and emerging challenges specific to demographic fairness in FR necessitate a dedicated review. In this article, we provide a consolidated discussion of recent advancements in the field, addressing the causes of demographic bias, available datasets for research, evaluation metrics for assessment of fairness, and recent mitigation techniques. Furthermore, we explore ongoing challenges that persist in addressing demographic disparities, particularly in the light of novel use cases and emerging FR applications. Although our primary focus is on race and ethnicity, as these are dominant areas of research, we also include gender-related studies within the broader context. Age-related studies, given their distinct nature and established body of research, are referenced only where directly relevant.
Naming Conventions: In this review, we use the term demographic bias to refer to accuracy differentials observed across demographic groups (e.g., race, gender, age). We adopt this terminology to remain consistent with much of the recent literature in FR and biometrics community. It may be noted that this term is not intended to imply intentional discrimination or moral judgment; rather, it reflects measurable differences in algorithmic performance across groupings that are often of regulatory and ethical interest.
@@
In this review, we adopt the terminologies used in the referenced works while acknowledging the minor differences in naming conventions. Terms such as race and ethnicity are often used interchangeably, although they represent distinct concepts. For clarity, we retain the terms employed by the original studies. Similarly, the names of ethnic groups vary across the literature: for instance, some works use terms like Black and White, while others prefer African (or African-American) and Caucasian. Additionally, the South Asian group is sometimes referred to as Indian, whereas Asians often refer to East Asians. To maintain consistency and respect the source material, we adhere to the original terminologies in this review. Readers are encouraged to refer to the cited works for precise definitions and context. While fairness in related topics such as face detection, image quality, expression recognition, and attribute estimation is of interest to the research community, this review exclusively focuses on demographic fairness in FR.
@@
Contributions: This review constitutes the first comprehensive work dedicated to exploring demographic fairness in FR, offering a unified and holistic perspective to researchers in the field. We systematically analyze and organize key aspects, including the causes of demographic disparities, available datasets, assessment metrics, and mitigation techniques, providing a structured framework for understanding these areas. Finally, we identify emerging challenges and unresolved questions, inviting further research and innovation to advance fairness in FR systems.
@@
The structure of this paper is as follows: Section II explores the causes of bias, analyzing factors such as distribution of demographic groups in datasets, skin-tone, image quality, and soft attributes. We provide an overview of datasets commonly used for fairness-related research in Section III, highlighting their demographic attributes and suitability for specific tasks. Section IV reviews existing metrics for fairness evaluation, discussing their strengths and limitations. Section V outlines recent bias mitigation strategies across different stages of the FR pipeline. We discuss open challenges and future research directions in Section VI, and conclude the review in Section VII.
II Causes of Demographic Differentials
@@
In this section, we consolidate findings from existing works related to causes of accuracy differentials in FR. Considering the wide range of research in this area, we have grouped the causes into categories for clarity and systematic review. These categories encompass factors such as imbalances in training datasets, variability in skin-tones, algorithmic sensitivity, image quality, related covariates, combined or intersectional factors, and soft attributes. Though this categorization simplifies the organization, it is important to note that many studies attribute accuracy differentials to multiple, overlapping factors, making strict classification difficult. We have categorized works based on their primary focus or findings, but some studies have been referenced across multiple categories to reflect their broader relevance. This approach also highlights that causes of demographic differentials or bias are inherently multifaceted and interconnected, requiring a thorough understanding to analyze and mitigate these disparities. A quick summary of various works discussed in this section can be found in Table I.
@@ II-A Training Datasets
Early studies in this area speculated that the issue of demographic bias in FR systems often stems from imbalanced or unrepresentative datasets, significantly influencing both training and evaluation outcomes. Research by Krishnapriya et al. [22] demonstrated how demographic groups, such as African-American cohorts, exhibit higher false match rates (FMR), while Caucasian cohorts face higher false non-match rates (FNMR), highlighting the interplay between race and matching thresholds. The Face Recognition Vendor Test (FRVT) conducted by NIST [15] substantiated these findings, reporting increased false positive identifications in women, children, and the elderly, alongside higher false negatives in under-represented racial groups, emphasizing the intricate interactions between dataset characteristics and demographic attributes.
@@
In a pre-deep learning era, studies like Klare et al.  [23] examined the impact of demographic balance in training datasets on FR performance. While their results indicated that underrepresented groups– such as females and Black individuals– benefited from increased representation, they also found that balancing the dataset alone did not fully eliminate disparities in recognition accuracy at test time. To further mitigate this, they proposed training exclusive cohort-specific models and using dynamic model selection during inference, highlighting that data balance might often be complemented with architectural or procedural adjustments to meaningfully reduce bias. Cavazos et al. [24] identified how dataset complexity and identification thresholds contribute to racial bias, such as the need for higher thresholds for East-Asian faces to achieve comparable false acceptance rates (FARs). Gwilliam et al. [25] challenged the prevailing assumptions about the necessity of balanced datasets by demonstrating that skewed distributions favoring African faces reduced racial differentials more effectively than balanced datasets. In [26, 27], Albiero et al. demonstrated that balancing training data for both genders (male and female) did not help reducing the corresponding differential in FR accuracy. Wu and Bowyer [28] expanded this discussion, emphasizing that mere balance in identities or number of images is insufficient to address disparities, highlighting additional factors like brightness and head pose during dataset assembly.
@@
Other works delved into specific aspects of demographic balance in datasets. Wang et al.  [29] observed that even race-balanced datasets failed to eliminate accuracy differentials, hypothesizing that certain ethnicities are inherently more challenging to recognize. Kolla and Savadamuthu [30] highlighted the influence of facial quality and racial feature gradations on model fairness. Focusing on inter-sectional bias, Muthukumar et al.  [31] identified structural facial features as significant contributors, particularly for dark-skinned females, over attributes like skin-tone or hair length. Cook et al. [32] further analyzed the role of image acquisition conditions, noting how factors such as skin reflectance and environmental conditions disproportionately affect darker-skinned individuals, thus advocating for standardized acquisition protocols to mitigate bias.
Although several studies acknowledge that significantly imbalanced training data contribute to unfair or biased FR models and training with balanced datasets reduces disparities to some extent, there is a consensus that these are neither the sole causes nor complete solutions to the broader issue of demographic fairness. Availability of large-scale training data, coupled with deeper models with better learning capacity, has made it possible to investigate the aspect of datasets in a more detailed manner. Recent findings have indicated that the role of balancedness of training data is limited, and other factors that are inherent to the individuals or their soft attributes may have a bigger role. Additionally, balancing training data (especially through sub-sampling majority groups) may involve discarding large amounts of high-quality data, thereby reducing overall training volume and potentially decreasing recognition accuracy while reducing the accuracy across constituent demographic groups.
@@ II-B Variability in Skin-tone
The influence of skin-tone on the performance of FR systems has been extensively studied, revealing significant demographic disparities. In [31], Muthukumar et al. identified notable under-performance in recognizing dark-skinned females compared to other demographic group for commercial classifiers. Their analysis attributed these disparities to structural features such as lips, eyes, and cheeks, in addition to skin-tone itself. In another study, Krishnapriya et al.  [22] examined FMR and FNMR across skin-tone groups, observing higher FMR for African-American cohorts and higher FNMR for Caucasians; however, they did not find a direct causation between darker skin-tones and higher error rates. Buolamwini and Gebru [33] employed the Fitzpatrick skin classification system to evaluate commercial gender classifiers and reported the lowest accuracy for darker-skinned females. While this study evaluated gender classification rather than identity recognition, it led to broader discussions around fairness in facial analysis systems, including those used in biometric applications.
@@
The Biometric Technology Rallies organized by MdTF have offered comprehensive insights into the role of skin-related factors in FR bias. Their 2019 report [32] emphasized skin reflectance as a more significant predictor of performance disparities than race. Using systematic linear modeling, their study demonstrated that darker skin-tones were associated with longer transaction (processing overall pipeline) times and lower accuracy in biometric systems. The longer transaction times were primarily attributed to difficulties in the face detection or image acquisition stage under suboptimal lighting conditions. Lower skin reflectance can reduce contrast, making it harder for detection algorithms to localize the face, thus increasing processing time. This dependency was found to vary substantially across systems, highlighting important role of acquisition methods in determining the extent of bias. Lu et al.  [34] provided a quantitative assessment of performance variations across five skin-tone groups, identifying light-skinned individuals as the easiest to verify and darker-skinned individuals as the most challenging. However, ambiguities in defining skin-tone categories complicate direct evaluations, highlighting the need for standardized classification metrics.
While many studies report that individuals with lighter skin-tones tend to be recognized more accurately than those with darker skin-tones, there is no consistent consensus that skin-tone is the primary driver of demographic bias. Moreover, it is important to note that most analyses rely on apparent or perceived skin tone–estimated via brightness or proxy classification schemes–rather than direct, objective measurement.
TABLE I: Summary of works delving into various causes of demographic fairness in face recognition. As several works have identified multiple causes of bias, we have categorized the works based on their primary focus or inference. For details, readers are encouraged to refer to the source materials. The demographic factors of primary interest are denoted as ET: Ethnicity or race, GN: gender or sex, AG: age; whereas + indicates study of more attributes.
@@
Continued on the next page.
TABLE II: (Continued)
Reference Year Dataset Attribute Summary
@@
Klare et al. [23] 2012 PCSO ET, GN, AG Designed dynamic face matchers to work with models trained on exclusive cohorts.
@@
Ricanek et al. [39] 2015 ITWCC AG Observed inefficiency in choice of algorithms towards recognizing children’s faces (due to structural changes).
@@
Pangelinan et al. [43] 2025 MORPH ET, GN, + Balancing brightness improves similarity scores for females.
@@
Bhatta et al.  [45] 2024 MORPH ET, GN Studied impact of blur and resolution for 1:N identification.
@@
Sarridis et al. [48] 2023 RFW ET, GN, AG Reported high error rates for African females over 60 years, highlighting compounded biases.
@@
Atzori et al. [50] 2022 VGGFace2, DiveFace ET, GN, + Explanatory framework for evaluating the influence of both protected and non-protected attributes .
@@
Albiero et al. [51] 2020 MORPH ET, AG Observed lack of consistent relation between recognition accuracy and age.
Albiero et al. [52] 2022 MORPH, MFAD Face region Studies impact of facial regions on bias patterns.
Bhatta et al. [53] 2023 MORPH, MFAD Hairstyle Noted that hairstyle balance reduces gender bias.
Pangelinan et al. [54] 2024 Various Face regions Higher FMR for females despite controlling other factors.
Ozturk et al. [55] 2024 MORPH Facial hair Study of facial hairs on impostor/genuine scores.
Wu et al. [56] 2024 MORPH, BA-test Facial hair Analysis of facial hairs (occlusions) for recognition accuracy.
Mamede et al. [57] 2024 RFW (occluded) Occlusions Reported that occlusions amplify racial bias.
Kurz et al. [58] 2025 MAAD-Face Facial Attributes Gender gap vanishes with matched attributes.
II-C Algorithmic Factors
In this section, we examine the sensitivity and limitations of algorithms in addressing demographic attributes. Phillips et al. [35] identified the “other-race effect,” where algorithms developed in Western and East Asian contexts demonstrated superior performance for their respective majority racial groups. This disparity persisted even when datasets were balanced, pointing to underlying biases in algorithmic design and training processes. Klare et al.  [23] observed recognition challenges for specific demographic groups, including females, Black individuals, and younger cohorts. They reported improved performance when models were trained exclusively on specific demographic cohorts. This improvement was supported by the integration of dynamic face matchers within the recognition pipeline—a modification that reflects both algorithmic adaptation and the influence of training cohort composition. In [36], Nagpal et al. demonstrated that deep learning models encode in-group biases, mirroring human tendencies such as own-race and own-age effects. By analyzing activation maps, they showed that these biases were ingrained within the feature representations of the models. Wang et al. [29], using the Racial Faces in-the-Wild (RFW) dataset, validated racial bias, revealing that error rates for African faces were nearly double those for Caucasians, even with race-balanced training data. Serna et al. [37] similarly highlighted significant performance gaps across demographic groups, advocating for diverse training datasets and fairness-aware algorithmic designs.
Further investigations into gender-based disparities by Albiero et al. [38] revealed skewed impostor and genuine score distributions as the primary reasons for lower accuracy in women. This disparity persisted across datasets, regardless of balanced training and neutral facial expressions. Ricanek et al. [39] noted unique challenges in recognizing children’s faces due to structural changes with age, finding that algorithms effective on adult faces performed poorly for younger subjects. Although this issue pertains to longitudinal variations in facial morphology, it underscores a broader implication: the same algorithms may not be equally effective across seemingly similar tasks (FR). Attributes such as facial structure should therefore be considered as critical factors in guiding the design and selection of appropriate recognition algorithms.
Identifying algorithmic factors that contribute to differential recognition accuracy across demographic groups remains a complex task. This process requires isolating the influence of the training algorithm–an integral component of the recognition pipeline. To ensure valid attribution, experimental setups need to be carefully controlled that differ in nothing but algorithms. Initially, performance disparities were primarily attributed to handful factors such as imbalances in training datasets or variations in acquisition conditions. However, as the number of potential influencing factors grows, the experimental studies to rigorously investigate and isolate the algorithmic sources of bias require a thorough and systematic approach.